# Articulatory Encodec: Vocal Tract Kinematics as a Codec for Speech

### v1

@@ III-B Articulatory Synthesis
We adopt HiFi-GAN as the vocoder for articulatory synthesis [56, 14, 44]. The vocoder is trained to synthesize speech audio with a 16K Hz sampling rate from articulatory features defined in §III-A1 and §III-A2. To condition the generation on the speaker embedding, we apply FiLM [57] to each convolution module with different receptive fields in the HiFi-GAN architecture, which modulates the output channels of each module. We adopt the same loss functions as [56]: Mel-spectrogram loss for reconstruction and multi-period and multi-scale discriminator loss for GAN training.
@@ III-C Dataset
For training the vocoder, we use LibriTTS-R [58], an enhanced version of LibriTTS [59]. The dataset is comprised of 585 hours of reading audiobooks ((555, 15, 15) hours for (train, dev, test) sets). The original 24K Hz audio is downsampled to 16K Hz. We use VCTK [60] to further evaluate the generalizability of the model to a broader range of speakers and accents. The entire VCTK dataset is not included in the training set and is only used for evaluation.
Note that only the FFN in the speaker identity encoding and the vocoder are updated in the large-scale training (orange modules in Fig. 2), and the rest of the pipeline remains fixed while training. More details of implementation and training are in Appendix B.1-7.
@@
IV Results & Experiments
We evaluate the proposed articulatory encodec in different aspects to validate each part of the framework – universality of articulatory encoding, IV-A, information conservation in encoding-decoding (or resynthesis) IV-B, and disentanglement of speaker embedding by speaker identity encoding IV-D.
IV-A Universality of Single-Speaker Articulatory Encoding
@@
As we claim a single-speaker AAI can be used for universal articulatory encoding, we compare our model with the existing state-of-the-art (SOTA) multi-speaker articulatory encoding systems that utilize tract variables [17, 28]. The tract variables are a set of articulatory parameters that can be derived from EMA and are known to be more regularized across speakers [61]. Table I denotes the performance of the tract-variable-based AAI systems and ours on two multi-speaker EMA datasets, MOCHA [62] and HPRC [63]. MOCHA includes 7 speakers with 27 minutes of data per speaker on average, and HPRC includes 8 speakers with 59 minutes of data per speaker on average. The performance is evaluated by the Pearson correlation coefficient (PCC), where results besides ours are retrieved from reference papers (95% confidence interval is denoted for our case.). Since our AAI is trained on a single-speaker (MNGU0) articulatory space, we fit a linear model that maps the MNGU0 articulatory space to another speaker’s articulatory space to measure the correlation [10]. As a result, our single-speaker AAI system shows correlations higher in MOCHA and slightly lower in HPRC compared to Wu et al. [17], and slightly higher in HPRC compared to Siriwardena and Espy-Wilson [28]. Note that this is not a fair comparison since the scores are measured on different variables. However, this system-wise comparison suggests that the single-speaker approach can yield a similar level of consistency across multiple speakers. Moreover, our approach only fits a linear model, thus it is likely to be more generalizable than the previous systems which use non-linear, deep recurrent neural networks. Fig. 3 demonstrates prediction examples of the predicted EMA. The left panel shows a near-perfect prediction even with a simple linear mapping from WavLM, and even after the affine transformation from MNGU0 space to other speakers, the predictions show high correlations with the ground truths (middle and right panel). The prediction performance on MNGU0 shows $0.878\pm 0.012$ average correlation (see Appendix A.2-4. for more analyses).
@@ IV-B Performance of Resynthesis by Articulatory Encodec
We evaluate the performance of the articulatory encodec in terms of intelligibility and quality of resynthesized speech audio. We used an out-of-the-box automatic speech recognition (ASR) model, Whisper (”openai/whisper-large-v3”), to evaluate the word error rate (WER) and character error rate (CER) of resynthesized speech. [^3] The transcription is normalized before measuring the error rates. The quality of the audio is measured by a human-evaluated subjective metric, mean opinion score (MOS), and a machine-evaluated MOS, UTMOS [64]. We evaluate both ground truth speech and resynthesized speech on the test-clean subset of LibriTTS-R which includes 8.56 hours from 39 unseen speakers. To further evaluate the generalizability of the model, we evaluate the model on VCTK data which includes 107 speakers with 500 utterances in total. Table II summarizes the performance in comparison with the ground truth speech. For the quality metrics, we reported 95% confidence intervals. When tested on LibriTTS-R, the ASR on the resynthesized speech shows high intelligibility with WER of 5.43% and CER of 2.90%, which is marginally different from those on the ground truth. Moreover, our articulatory synthesis generates natural speech sounds showing decent quality with MOS of 3.82 which is a 0.19 decrease from the ground truth, and with UTOMOS of 4.12 which is on par with the ground truth.
@@
Furthermore, the results on VCTK also demonstrate high intelligibility with WER of 3.73% and CER of 1.96%. Though the gap from the ground truth is larger than LibriTTS-R, this is a remarkable level of performance given that the model is not exposed to the entire dataset during training. As the training set audio corpus which has cleaner audio than VCTK, the model demonstrates some speech enhancement capacity which results in a higher MOS and UTMOS than the ground truth. While the speakers in LibriTTS-R are concentrated on American English, VCTK includes a variety of accents, primarily focusing on English speakers with different regional accents from the United Kingdom. Therefore, the high performance demonstrated on VCTK suggests that the articulatory encodec can robustly generalize to unseen speakers and accents.
IV-C Multilingual Application
TABLE III: Intelligibility of multilingual application. WER and CER on ground truth speech (GT), resynthesized speech by English-only trained model (EN-R), and by multilingual model (ML-R).
@@ IV-C Multilingual Application
GT EN-R ML-R GT EN-R ML-R
German 5.23 15.32 9.70 1.89 6.66 3.95
Dutch 9.90 20.20 15.39 2.82 8.83 5.70
Portuguese 7.10 17.81 12.80 2.19 8.50 5.21
Italian 9.63 22.09 15.53 2.43 7.40 4.98
Polish 4.19 14.79 9.89 0.94 4.63 2.87
Spanish 3.96 10.74 6.79 1.26 4.73 2.38
French 4.62 16.67 8.04 2.65 9.90 4.54
Korean 6.31 15.88 13.66 1.01 6.72 4.96
Japanese 5.65 8.45 8.16
Chinese 10.69 31.27 23.28
Average 6.37 16.69 11.48 3.15 9.71 6.60
@@
We evaluate the articulatory encodec on speech corpora from other languages: 7 European languages (German, Dutch, Portuguese, Italian, Polish, Spanish, French) from multilingual LibriSpeech [65] and 3 East Asian languages (Korean, Japanese, and Chinese (Mandarin)) from KSS [66], JVS [67], and AISHELL [68], respectively. Except for KSS, all corpora include multiple speakers. We randomly sampled 200 utterances from the test split of each corpus to evaluate the intelligibility of resynthesized speech. [^4] We excluded samples with numbers to avoid converting Arabic numerals. Table III compares WER and CER by ASR (Whisper) in each language. As no space is used for Japanese and Chinese, we omit the WER for these languages. We also evaluate the performance of a version of the model that is fine-tuned to the multilingual data, which is denoted as ML-R in Table III). The original English-only trained encodec is denoted as EN-R. See Appendix B.8. for details of the fine-tuning procedure.
As shown in Table III, with only training in English, the resynthesis by articulatory encodec achieves average WER of 16.69 excluding Japanese and Chinese, average CER of 9.71. Roughly, if the conflated Chinese CER is excluded, the EN-R preserves a fair amount of information given that 93% of characters are correctly recognizable. Despite huge gaps from the ground truth observed, this is remarkable generalizability given both encoding and decoding procedures have only seen English in training. When fine-tuned, the average WER and CER are cut down to 11.48% and 6.60%, respectively. This indicates that some portion of the errors is induced by the out-of-domain application of the vocoder, thus fine-tuning is able to yield a huge gain in performance. Yet, there are unresolved gaps from the ground truth, which would be attributed to the fact that we are only tuning the vocoder and keeping the encoder part intact. This is also aligned with the observation that there is a slight language bias in the articulatory representation in SSL [10].
@@ IV-D Speaker Recognition and Voice Control by Speaker Identity
The speaker identity encoding is especially important to inform voice texture for speech synthesis since the proposed articulatory encoding is agnostic to speakers. On one hand, this suggests that the model can learn a speaker embedding that is disentangled from individual articulatory patterns, or dialects and accents. Here, we evaluate this claim by experiments on few-shot speaker identification and zero-shot voice conversion.
First, we build a few-shot, learning-free speaker identification (SID) by comparing the similarity between the speaker embeddings. For each test speaker, 10 first clips in the dataset are concatenated and then the template speaker embedding for the speaker is extracted. Then among the potential test speakers, the speaker is identified as the speaker with maximal similarity in the template and tested speaker embedding. For LibriTTS-R, we use speakers in the test-clean set with at least 10 clips, leaving 36 speakers. For VCTK, we use the train set clips to create the template embeddings for 107 test speakers. Additionally, we also evaluate SID accuracy using all data in the train set for creating the templates, which is denoted in parentheses in Table IV.
@@
We further evaluate the validity of our identity encoding by zero-shot voice cloning application. We selected 6 speakers from VCTK, where their mean pitch values are equidistant in the range from 80 Hz to 230 Hz. Fig. 4 visualizes the speaker embedding of the selected speakers, demonstrating that each speaker’s utterances are clustered distinctively. On the contrary, no speaker-relevant cluster is identifiable in the manifold of utterance-wise averaged articulations, indicating disentanglement between articulation and speaker identity. For each target speaker, the first 10 clips in the training set are concatenated and then the speaker embedding is extracted. Then, the voice from the source speaker is converted by switching speaker embedding in the resynthesis. Additionally, the pitch range of the source speaker is adjusted to match that of the target speaker, by z-scoring the source pitch trace; and shifting and rescaling to match the center and scale of the target pitch range (”P-rescale”). We convert all audio clips in test-clean set of LibriTTS-R with more than 2 seconds of duration, to each of 6 VCTK target speakers.
We evaluate “coding-recoding similarity” of the voice-converted speech. The coding-recoding similarity measures the correlation between the original input speech coding and the encoding of the output of the articulatory synthesis, where the voice is controlled. For articulation, we measure the Pearson correlation of each of 12 channels and then average them, and for pitch and loudness, the correlation is separately reported. For speaker identity, the cosine similarity between the conditioning target embedding and that of the generated speech. Furthermore, we report the SID accuracy based on the task designed in the above few-shot SID, and WER and UTMOS to check the intelligibility and quality of the converted audio. As baselines, we evaluate existing voice conversion models: YourTTS [74], a multi-speaker TTS that offers zero-shot voice conversion, and QuickVC [47], an SSL-based encodec model which conditions voice on an acoustics-based speaker encoding. The former is included as an example of a voice conversion with text-prior, and the latter is included since the model has a similar SSL-based encoding but without the articulatory constraint. [^7] Up to our knowledge, this is the only open-sourced SSL-based encodec with an official checkpoint. Furthermore, we evaluate the coding-recoding similarity of the self-targeted resynthesized speech (“Ours-Resynth” in Table. V). The results are summarized in Table. V with 95% confidence intervals if applicable.
TABLE V: Performance comparison of Zero-Shot Voice conversion
@@
YourTTS VCTK $0.923\pm 0.021$ $0.881\pm 0.165$ $0.881\pm 0.062$ $0.817\pm 0.080$ 81.4 4.00 $4.05\pm 0.51$
QuickVC VCTK $0.930\pm 0.021$ $0.868\pm 0.172$ $0.903\pm 0.063$ $0.831\pm 0.070$ 66.7 4.17 $4.22\pm 0.36$
@@
Ours-Resynth Self $0.954\pm 0.014$ $0.968\pm 0.109$ $0.936\pm 0.077$ $0.962\pm 0.038$ 91.5 4.40 $4.00\pm 0.59$
As a result, the voice-converted speech by the articulatory encodec can be accurately identified as the target speaker among 107 VCTK speakers, achieving 94.6 % SID accuracy (Table. V). Though the speaker coding-recoding similarity is lower than the self-targeted resynthesis (0.879 and 0.962, respectively), the SID result suggests that this level of similarity is high enough to be discriminable from more than a hundred unseen speakers. Furthermore, our approach shows higher SID accuracy than both of the baseline models, where the speaker coding-recoding similarities are also inferior to ours.
@@
The high coding-recoding similarity in articulation suggests that the articulation is well disentangled from speaker identity and pitch, which indicates that the dialect is preserved after the voice conversion. The correlation is as high as 0.944 and is marginally different from the self-targeted resynthesis, 0.954, which also indicates the stability of the coding by our articulatory encodec. This result further supports that the proposed AAI is agnostic to speaker identity. Moreover, the similarities outperform baselines, especially in pitch, suggesting our model better disentangles the pitch. Lastly, the intelligibility slightly decreases after the voice conversion, but the difference is marginal and the speech remains highly intelligible with WER of 4.83 %. The YourTTS converted speech shows a lower WER of 4.00 %, which might be driven by the text-to-speech (TTS) prior in the model. Finally, the voice conversion shows some degraded UTMOS as 3.83, 0.17 decrease compared to the resynthesis, and QuickVC shows the highest UTMOS as 4.22. This suggests that our model has a potential trade-off between disentanglement and audio quality of the synthesized audio. However, this degrades in naturalness may be confounded by a natural correlation between accents and voice textures, so that some voice conversion cases may result in rare combinations of accent and speaker identity, harming the UTMOS score. To sum up, the articulatory encodec is a stable and disentangled coding system of speech.

### v2

@@ III-B Articulatory Synthesis
We adopt HiFi-GAN as the vocoder for articulatory synthesis [60, 14, 47]. The vocoder is trained to synthesize speech audio with a 16K Hz sampling rate from the articulatory features. To condition on the speaker embedding, we apply FiLM [61] to each convolution module in the HiFi-GAN architecture, which modulates the output channels of each module. We adopt the same loss functions as [60]: mel spectrogram loss for reconstruction and multi-period and multi-scale discriminator loss for GAN training.
@@ III-C Dataset
For training the vocoder and speaker encoder, we use LibriTTS-R [62], an enhanced version of LibriTTS [63]. The dataset is comprised of 585 hours of reading audiobooks (555 hours for training). The original 24K Hz audio is downsampled to 16K Hz. We use VCTK [64] to further evaluate the generalizability of the model to a broader range of speakers and accents. The entire VCTK dataset is unseen during training and only used for evaluation.
Note that the FFN in the speaker identity encoding and the HiFi-GAN vocoder are the only trainable modules (orange modules in Fig. 2), and the rest of the pipeline remains fixed while training. More details of implementation and training are in Appendix B.1-7.
IV Experiments & Results
We evaluate the proposed articulatory encodec in different aspects to validate each part of the framework – universality of articulatory encoding (IV-A), information conservation in encoding-decoding (IV-B), multilingual generalization (IV-C), and disentanglement of speaker identity encoding (IV-D).
IV-A Universality of Single-Speaker Articulatory Encoding
As we claim a single-speaker AAI can be used for universal articulatory encoding, we compare our model with the existing state-of-the-art (SOTA) multi-speaker articulatory encoding systems that utilize tract variables [17, 28]. The tract variables are a set of articulatory parameters that can be derived from EMA and are known to be more regularized across speakers [65]. Table I denotes the performance of the tract-variable-based AAI systems and ours on two multi-speaker EMA datasets, MOCHA [66] and HPRC [67]. MOCHA includes 7 speakers with 27 minutes of data per speaker on average, and HPRC includes 8 speakers with 59 minutes of data per speaker on average. The performance is evaluated by the Pearson correlation coefficient (PCC), where results besides ours are retrieved from reference papers. (95% confidence interval is denoted for our case.) Since our AAI is trained on a single-speaker (MNGU0) articulatory space, the model outputs are not directly comparable to the EMA from MOCHA-TIMIT or HPRC. Therefore, we fit a linear model to spatially align MNGU0’s articulatory space to another speaker’s articulatory space to measure the correlation [10]. (See Appendix A.2. for details.) As a result, our single-speaker AAI system shows correlations higher in MOCHA and slightly lower in HPRC compared to Wu et al. [17], and slightly higher in HPRC compared to Siriwardena and Espy-Wilson [28]. Note that this is not a fair comparison since the scores are measured on different variables. However, this system-wise comparison suggests that the single-speaker approach can yield a similar level of consistency across multiple speakers. Moreover, our approach only fits a linear model, thus it is likely to be more generalizable than the previous systems which use non-linear, deep recurrent neural networks. Fig. 3 demonstrates prediction examples of our AAI model. The left panel shows a near-perfect prediction even with a simple linear mapping from WavLM, and even after the affine transformation from MNGU0 space to other speakers, the predictions show high correlations with the ground truths (middle and right panel). The prediction performance on MNGU0 shows $0.878\pm 0.012$ average correlation, and additional analyses on AAI are reported in Appendix A.3-5.
@@ IV-B Performance of Resynthesis by Articulatory Encodec
We measure the intelligibility and quality of resynthesized speech audio to evaluate how well information is preserved in the encoding-decoding process. We used an out-of-the-box automatic speech recognition (ASR) model, Whisper (“openai/whisper-large-v3”) [68], to evaluate the word error rate (WER) and character error rate (CER) of resynthesized speech. [^3] The transcription is normalized before measuring the error rates. The quality of the audio is measured by a human-evaluated subjective metric, mean opinion score (MOS), and a machine-evaluated MOS, UTMOS [69]. We evaluate both ground truth speech and resynthesized speech on the test-clean subset of LibriTTS-R which includes 8.56 hours from 39 unseen speakers. To further evaluate the generalizability of the model, we evaluate the model on VCTK dataset that includes 107 speakers with 500 utterances in total. Table II summarizes the performance in comparison with the ground truth speech. For the quality metrics, we reported 95% confidence intervals. When tested on LibriTTS-R, the ASR on the resynthesized speech shows high intelligibility with WER of 5.43% and CER of 2.90%, which is marginally different from those of the ground truth. Moreover, our articulatory synthesis generates natural speech sounds showing decent quality with MOS of 3.82 which is a 0.19 decrease from the ground truth, and with UTOMOS of 4.12 which is on par with the ground truth.
Furthermore, the results on VCTK also demonstrate high intelligibility with WER of 3.73% and CER of 1.96%. Though the gap from the ground truth is larger than LibriTTS-R, this is a remarkable level of performance given that the model is not exposed to the entire dataset during training. As LibriTTS-R has cleaner audio than VCTK, the model demonstrates some speech enhancement capacity which results in a higher MOS and UTMOS than the ground truth. While the speakers in LibriTTS-R are concentrated on American English, VCTK includes a variety of accents, primarily focusing on English speakers with different regional accents from the United Kingdom. Therefore, the high performance demonstrated on VCTK suggests that Articulatory Encodec can robustly generalize to unseen speakers and accents.
IV-C Multilingual Generalization
TABLE III: Performance comparison on multilingual speech data. Each metric is evaluated on ground truth speech (GT) and resynthesized speech by English-only trained model (EN-R) or multilingual model (ML-R).
@@ IV-C Multilingual Generalization
GT EN-R ML-R GT EN-R ML-R GT EN-R ML-R
German 5.23 15.32 9.70 1.89 6.66 3.95 $2.96\pm 0.83$ $2.87\pm 0.58$ $2.83\pm 0.72$
Dutch 9.90 20.20 15.39 2.82 8.83 5.70 $2.91\pm 0.82$ $2.89\pm 0.65$ $2.77\pm 0.79$
Portuguese 7.10 17.81 12.80 2.19 8.50 5.21 $2.63\pm 1.34$ $2.81\pm 0.67$ $2.56\pm 0.92$
Italian 9.63 22.09 15.53 2.43 7.40 4.98 $2.78\pm 1.26$ $2.82\pm 0.61$ $2.67\pm 0.82$
Polish 4.19 14.79 9.89 0.94 4.63 2.87 $3.09\pm 0.76$ $3.00\pm 0.63$ $2.95\pm 0.76$
Spanish 3.96 10.74 6.79 1.26 4.73 2.38 $2.93\pm 0.86$ $2.88\pm 0.66$ $2.72\pm 0.81$
French 4.62 16.67 8.04 2.65 9.90 4.54 $2.79\pm 0.93$ $2.80\pm 0.65$ $2.69\pm 0.79$
Korean 6.31 15.88 13.66 1.01 6.72 4.96 $3.69\pm 0.63$ $3.47\pm 0.80$ $2.91\pm 0.95$
Japanese 5.65 8.45 8.16 $3.53\pm 0.80$ $3.49\pm 0.77$ $3.08\pm 0.83$
Chinese 10.69 31.27 23.28 $2.66\pm 1.21$ $3.06\pm 0.92$ $2.71\pm 1.12$
Average 6.37 16.69 11.48 3.15 9.71 6.60 $3.00\pm 0.66$ $3.01\pm 0.49$ $2.79\pm 0.29$
@@
We evaluate Articulatory Encodec on speech corpora from other languages: 7 European languages (German, Dutch, Portuguese, Italian, Polish, Spanish, French) from multilingual LibriSpeech [70] and 3 East Asian languages (Korean, Japanese, and Chinese (Mandarin)) from KSS [71], JVS [72], and AISHELL [73], respectively. Except for KSS, all corpora include multiple speakers. We randomly sampled 200 utterances from the test split of each corpus to evaluate the resynthesized speech. [^4] We excluded samples with numerals to avoid converting Arabic numerals. Table III compares the intelligibility (WER and CER using Whisper) and quality (UTMOS). As no interword spacing is used for Japanese and Chinese orthographic systems, words in these languages are not separated as straightforwardly as other languages. Therefore, we omit the WER for these languages. We also evaluate the performance of a multilingual version of the model that is obtained by fine-tuning to the multilingual data (denoted as ML-R in Table III). The original English-only trained encodec is denoted as EN-R. See Appendix B.8. for details of the fine-tuning procedure.
As shown in Table III, the resynthesis by English-only model (EN-R) achieves average WER of 16.69 excluding Japanese and Chinese, average CER of 9.71. Roughly, if the conflated Chinese CER is excluded, the EN-R preserves a fair amount of information given that 93% of characters are correctly recognized. Despite the huge gap from the ground truth, this is remarkable generalizability given both encoding and decoding procedures have only seen English in training. When fine-tuned (ML-R), the average WER and CER are cut down to 11.48% and 6.60%, respectively. This indicates that some portion of the errors is induced by the out-of-domain application of the vocoder, thus that fine-tuning is able to yield a huge gain in performance. Yet, there are unresolved gaps from the ground truth, which would be attributed to the fact that we are only tuning the vocoder and keeping the encoder part intact. This is also aligned with the observation that there is a slight language bias in the articulatory representation in SSL [10]. In terms of synthesis quality, we find a minimal difference between ground truth ($3.00\pm 0.66$) and resynthesized audio from English-only trained model ($3.00\pm 0.66$), and the quality slightly drops to $2.79\pm 0.29$ when the model is fine-tuned to multilingual dataset. (The scores are averaged across languages and the ranges denote 95% confidence intervals.) The drop in quality is likely due to the general inferiority in the quality of multilingual datasets compared to LibriTTS-R.
@@ IV-D Speaker Recognition and Voice Control by Speaker Identity
The speaker identity encoding informs voice texture of speech, which is especially important as our articulatory features are speaker agnostic. On one hand, this suggests that the model can learn a speaker embedding that is disentangled from individual tendencies in articulation, or accents. Here, we evaluate this claim by experiments on few-shot speaker identification and zero-shot voice conversion.
First, we build a few-shot, learning-free speaker identification (SID) by comparing the similarity between the speaker embeddings. For each test speaker, 10 first clips in the dataset are concatenated and then the speaker embedding is extracted to serve as a template embedding for the speaker. Then among potential test speakers, the identity is predicted by choosing the speaker with maximal similarity between the template and tested speaker embedding. For LibriTTS-R, we use speakers in the test-clean set with at least 10 clips, leaving 36 speakers. For VCTK, we use the train set clips to create the template embeddings for 107 test speakers. Additionally, we also evaluate SID accuracy using all data in the train set for creating the templates, which is denoted in parentheses in Table IV.
@@
We further evaluate the validity of our identity encoding by zero-shot voice cloning application. We selected 6 speakers from VCTK, where their mean pitch values are equidistant in the range from 80 Hz to 230 Hz. Fig. 4 visualizes the speaker embedding of the selected speakers, demonstrating that each speaker’s utterances are clustered distinctively. On the contrary, no speaker-relevant cluster is identifiable in the manifold of utterance-wise averaged articulations. For each target speaker, the first 10 clips in the training set are concatenated and then the speaker embedding is extracted. Then, the voice from the source speaker is converted by switching speaker embedding in the synthesis. Additionally, the pitch range of the source speaker is adjusted to match that of the target speaker, by z-scoring the source pitch trace; and shifting and rescaling to match the center and scale of the target pitch range (”P-rescale”). We convert all audio clips in test-clean set of LibriTTS-R with more than 2 seconds of duration, to each of 6 VCTK target speakers.
TABLE V: Performance comparison of Zero-Shot Voice conversion. The SID accuracy using our speaker encoding is denoted in parenthesis.
@@
FreeVC VCTK $0.923\pm 0.021$ $0.881\pm 0.165$ $0.881\pm 0.062$ $0.839\pm 0.109$ 92.5 (81.4) 4.00 $4.05\pm 0.51$
QuickVC VCTK $0.930\pm 0.021$ $0.868\pm 0.172$ $0.903\pm 0.063$ $0.971\pm 0.105$ 65.0 (66.7) 4.17 $4.22\pm 0.36$
@@
Ours-Resynth Self $0.954\pm 0.014$ $0.968\pm 0.109$ $0.936\pm 0.077$ $0.962\pm 0.038$ 4.40 $4.00\pm 0.59$
We evaluate “coding-recoding similarity” of the voice-converted speech. The coding-recoding similarity measures the correlation between the speech coding of original speech and that of synthesized speech. Note that our coding pipeline is also an analytic measurement that provides articulatory and source features of speech. Therefore, the metric indicates how much the articulation and source information is preserved in the voice conversion and speech synthesis process. For articulation, we measure the Pearson correlation of each of 12 channels and then average them, and for pitch and loudness, the correlation is separately reported. For speaker identity, the cosine similarity is measured between the conditioning target embedding and the embedding extracted from the generated speech. Unlike articulatory and source features, speaker embedding is arbitrary and not interpretable. Therefore, we evaluate the accuracy in SID task, the same discriminability task designed above for VCTK. As the task requires to pinpoint the target speaker form 107 speakers, the SID accuracy can inform how close the voice converted speech to the target speaker. Lastly, WER and UTMOS to check the intelligibility and quality of the converted audio. As baselines, we evaluate existing voice conversion models: FreeVC [79] and QuickVC [50]. Both models encode speech contents using speech SSL models and have separate speaker encoders jointly trained with synthesizers. The former is more comparable to ours as the content encoder is also based on WavLM. None of the baseline models impose any grounded nor interpretable structures on the speech codes. We use our articulatory encoding to analyze articulatory and source features. The correlations of speaker embeddings and SID scores are evaluated using their own speaker encoders, where we also denote the SID accuracies using our speaker encoder in parentheses. Furthermore, we evaluate the coding-recoding similarity of the self-targeted resynthesized speech (“Ours-Resynth” in Table. V), to demonstrate consistency of the speech contents in synthesis without voice conversion. To avoid confusion, we do not report SID for resynthesis as the speaker identity is not imposed to be VCTK speakers. The results are summarized in Table. V with 95% confidence intervals if applicable.
As a result, the voice-converted speech by our model can be accurately identified as the target speaker among 107 VCTK speakers, achieving 94.6 % accuracy (Table. V). Though the similarity in speaker embedding is lower in voice conversion than in the self-targeted resynthesis (0.879 and 0.962, respectively), the SID result suggests that this level of similarity is high enough to be discriminable from 107 unseen speakers. Furthermore, our approach shows higher SID accuracy than both of the baseline models. FreeVC shows inferior scores in both speaker similarity and SID scores. QuickVC shows higher speaker similarity than ours but shows significantly lower SID accruacy as 65% (66.7% using our speaker encoding). The latter case indicates that their speaker embedding is relatively simple to be more consistent through voice conversion, but lacks specificity to discriminate the target speaker.
@@
The high coding-recoding similarity in articulation suggests that the articulation is well disentangled from speaker identity and pitch, which indicates that accent is preserved after the voice conversion. The correlation is as high as 0.944 and is marginally different from the self-targeted resynthesis, 0.954, which also indicates the stability of the coding by our articulatory encodec. This result further supports the claim that the proposed AAI is agnostic to speaker identity, yielding highly consistent inferences even with different voice identities. Moreover, the similarities outperform baselines, especially in pitch, suggesting our model better disentangles pitch from the other features. Lastly, the intelligibility slightly decreases after the voice conversion, but the difference is marginal and the speech remains highly intelligible with WER of 4.83 %. Both baselines shows lower WERs than our model (FreeVC: 4.00 %, and QuickVC: 4.17 %0) indicating a potential internal mechanism that might improve intelligibility by modifying the articulatory features of the original speech. Finally, the voice conversion shows some degraded UTMOS as 3.83, 0.17 decrease compared to the resynthesis, and QuickVC shows the highest UTMOS as 4.22. This suggests that our model has a potential trade-off between disentanglement and audio quality of the synthesized audio. However, this degrades in naturalness may be confounded by a natural correlation between accents and voice textures, so that some voice conversion cases may result in rare combinations of accent and speaker identity, harming the UTMOS score. To sum up, these results indicate that Articulatory Encodec is a stable and disentangled coding system of speech.